# Government Transparency Affects Innovation: Evidence from Wireless Products

### v1

@@ 6 Empirical strategy
I evaluate the effect of transparency on follow-on innovation on a subset of products that introduced a previously unused frequency combination (NewFrequency = 1). I estimate the causal effect of a product application having exhibits published online, Post, on the number of products that use the new frequency combination in the next five years (ForwardUse).
Simply regressing Post on ForwardUse would be insufficient for this purpose. Selection into the transparency treatment in the uptake period is voluntary and could bias the estimates. For example, more innovative firms might more likely introduce new frequencies and voluntarily use the digitalized submission process; or they could try to avoid disclosure associated with digitalization. A time-dependent confounder could affect the outcome in the pre- or post-period and bias the results. To overcome these challenges, I use two causal identification strategies: difference-in-differences and regression discontinuity designs. The difference-in-differences design enables estimating the change over time and on sub-populations. The regression discontinuity design enables estimating effects on alternative outcomes.
@@ 6.1 Difference-in-differences
$\bar{Product_{i}}$ denotes a vector of product random effects which I add to absorb the effects of product-level characteristics and time-determined policy changes. I use random effects instead of fixed effects because of the limited number of observations per unit, the over-dispersion of the outcome variable, and the convergence issues with fixed effects models. The random effects partially pool information across products to generate stable conservative coefficient estimates (Clark and Linzer, 2015; Bell et al., 2019).
As the outcome variable is an over-dispersed count, I use a negative-binomial model. I cluster standard errors on the frequency level. 88endnote: 8In a rare case, more than one product is introduced on the same date with the same new frequency combinations. I keep both products. As I expect the effect to differ in time, I present results from observations two, four, and ten years post the uptake period, with four years in the pre-period. I also estimate the dynamic treatment effect, where I interact the treatment group with year-fixed effects. Finally, I include a binary-outcome logistic regression and non-zero observations negative binomial regression with fixed effects as a robustness check.
The causal identification rests on the assumption of parallel trends: absent the transparency shock, the forward use of newly introduced frequency combinations would evolve similarly for the applicant and competitors. I cannot directly test this assumption. However, linear pre-trends in Table LABEL:apx:tab:diff_time (Wing et al., 2018) and staggered regression plots in Figure 4 do not show significant pre-trends.
@@ 6.2 Discontinuity
I use the sharp regression discontinuity design, which fits a non-parametric local regression following Calonico et al. (2015). I rely on robust estimates and report results from polynomials of degree one (main), two, and three, and let the algorithm choose optimal bandwidth from ten years of observations prior and post the shock, and use a triangular kernel.
The uptake period complicates the discontinuity design (Noack and Rothe, 2023). Excluding observations from the uptake period (so-called donut model) risks introducing artificial discontinuity and noise. Including observations from the uptake period risks underestimating the treatment effects, because treatment was incomplete. Thanks to the fast uptake (over 90 percent in April 1998), this downward bias might be limited. I present results from both approaches, which are broadly consistent. I refer to estimates from the specifications inclusive of the treatment period as the main (conservative) results.
@@ 7.2 Foreign and domestic competition
Foreign, Domestic $\times$ AIPA Yes Yes Yes Yes
@@ 8 Conclusion
Government regulations affect the information environment in which firms innovate. To elucidate this dynamic, I test the causal effects of increased regulatory transparency on follow-on innovation. I focus on a late-1990s transparency shock that decreased the costs of accessing detailed technical information about all wireless-enabled devices marketed in the U.S., a globally representative high-tech industry.
The results indicate that government transparency increased business dynamism, as evidenced by increased market entry and exit, and facilitated the adoption of new technologies. The estimated negative effect on the survival of the original firm exposes a familiar tension in innovation policy: eased access to a new technology benefits competitors and consumers but may hinder the original firm’s ability to offset the costs of developing the technology in the first place (Hall et al., 2014). In the long term, this dynamic may discourage innovation. These considerations preclude an unequivocally positive evaluation of increased transparency’s overall effects on innovation, as such an evaluation necessitates modeling firms’ strategic responses within a dynamic equilibrium of the altered information environment (Hall and Sena, 2017; Bryan and Williams, 2021).
The main message from my analysis is that government transparency can have large effects on business innovation in ways that are typically not at the forefront of policy debates on transparency. As of this writing, the FCC product database remains unaffected by the discontinuations, although numerous other government websites have been terminated (Gotfredsen, 2025). The findings highlight the information frictions present in innovation markets, and the power of government interventions in shaping the information environment in which firms innovate. Even regulations not directly aimed at supporting innovation can generate significant unintended effects by influencing information costs. As governments consider policies affecting information availability—such as algorithmic transparency, AI regulations, or restrictions on non-compete agreements—these results suggest potential secondary effects on innovation, even when this is not the primary objective.

### v2

@@ 6 Empirical strategy
I evaluate the effect of transparency on follow-on innovation on a subset of products that introduced a previously unused frequency combination (NewFrequency = 1). I estimate the effect of a product application having exhibits published online, Post, on the number of products that use the new frequency combination in the next five years (ForwardUse).
Simply regressing Post on ForwardUse would be insufficient for this purpose. Selection into the transparency treatment in the uptake period is voluntary and could bias the estimates. For example, more innovative firms might more likely introduce new frequencies and voluntarily use the digitalized submission process; or they could try to avoid disclosure associated with digitalization. A time-dependent confounder could affect the outcome in the pre- or post-period and bias the results. To overcome these challenges, I use two strategies: difference-in-differences and regression discontinuity designs. The difference-in-differences design enables estimating the change over time and on sub-populations. The regression discontinuity design enables estimating effects on alternative outcomes.
@@ 6.1 Difference-in-differences
$\bar{Product_{i}}$ denotes a vector of product random effects which I add to absorb the effects of product-level characteristics and time-determined policy changes. I use random effects instead of fixed effects because of the limited number of observations per unit, the over-dispersion of the outcome variable, and the convergence issues with fixed effects models. The random effects partially pool information across products to generate stable conservative coefficient estimates (clark_should_2015; bell_fixed_2019).
As the outcome variable is an over-dispersed count, I use a negative-binomial model. I cluster standard errors on the frequency level.  [^8] In a rare case, more than one product is introduced on the same date with the same new frequency combinations. I keep both products. As I expect the effect to differ in time, I present results from observations two, four, and ten years post the uptake period, with four years in the pre-period. I also estimate the dynamic treatment effect, where I interact the treatment group with year-fixed effects. Finally, I include a binary-outcome logistic regression and non-zero observations negative binomial regression with fixed effects as a robustness check.
The model rests on the assumption of parallel trends: absent the transparency shock, the forward use of newly introduced frequency combinations would evolve similarly for the applicant and competitors. I cannot directly test this assumption. However, linear pre-trends in Table LABEL:apx:tab:diff_time (wing_designing_2018) and staggered regression plots in Figure 4 do not show significant pre-trends.
@@ 6.2 Discontinuity
I use the sharp regression discontinuity design, which fits a non-parametric local regression following calonico_rdrobust_2015. I rely on robust estimates and report results from polynomials of degree one (main), two, and three, and let the algorithm choose optimal bandwidth from ten years of observations prior and post the shock, and use a triangular kernel.
The uptake period complicates the discontinuity design (noack_donut_2023). Excluding observations from the uptake period (so-called donut model) risks introducing artificial discontinuity and noise. Including observations from the uptake period risks underestimating the treatment effects, because treatment was incomplete. Thanks to the fast uptake (over 90 percent in April 1998), this downward bias might be limited. I present results from both approaches, which are broadly consistent. I refer to estimates from the specifications inclusive of the treatment period as the main (conservative) results.
@@
6.3 Empirical context
Several other changes occurred during the relevant time period that contextualize the transparency shock, and risk conflating the estimates of its effects. The founding of the WTO in 1995 and the accession of various countries in the subsequent decade establishes a potential confounder. Trade liberalization may have increased follow-on product introductions by both originators and competitors, altered the share of originator products by non-U.S. firms, or raised follow-ons to U.S. products—none of which are time-invariant. To a large degree, the difference-in-differences design addresses these concerns by comparing forward use by competitors (treatment) to forward use by the originator (control), which should be similarly exposed to aggregate, time-varying shocks. The regression discontinuity design puts the most weight on observations near the cutoff date. The cutoff lays two to three years after most countries joined the WTO in 1995 and 1996, and three to four years prior to the accession of China and Taiwan in 2001 and 2002.
A remaining threat is differential exposure when originators and competitors are located in countries affected differently by trade liberalization. To elucidate these cross-country patterns, table 4 reports sub-population estimates by competitors in the same country as the originator—including a U.S.-only subset (column 3), where trade liberalization should not differentially affect treatment vs. control; U.S. competitors following-on non-U.S. originators (column 4), where liberalization would bias effects downward; and splits by early (pre-1998) WTO membership and separate indicators for China and Taiwan (columns 2–3). If 1995–1996 WTO entry drove the results, we would observe breakpoints in the pre-period. Instead, figure 4 shows flat pre-trends over the three years preceding the transparency shock—aligning with the 1995 WTO founding. China and Taiwan are two major countries joining after the transparency shock, with significant imports to the U.S. Table LABEL:apx:tab:diff_no_china_taiwan presents main results excluding firms from China and Taiwan from the analysis, both as originators and as competitors. All these results, discussed in detail later, show significant, if reduced, effect on follow-on product introduction by domestic and early WTO members, providing evidence that the results are not driven by trade liberalization.
Domestically, the Electronic Freedom of Information Act of 1996 mandated government agencies to publish specific records electronically by the end of 1999 (usc_electronic_1996). I did not find any changes in FCC policies beyond the transparency changes detailed above. The Telecommunications Act of 1996 significantly revised the original 1934 law, focusing on enabling market entry in internet telecommunications services markets and increasing interoperability and standardization. It did not directly change the technical requirements in the product authorization process, but an amendment to Section 302 enabled the delegation of equipment testing and certification to private laboratories (usc_telecommunications_1996; emeritz_telecommunications_1996). This meant that the mandatory testing documentation firms submit in their authorization applications became more often provided by an external testing provider. The strategies outlined below are robust to the effects of these changes, as products applied for in the treatment and control groups and in the pre- and post-periods would be affected equally by this change.
The American Inventor’s Protection Act (AIPA) of 1999 accelerated disclosure of patent applications. Prior literature found a positive effect on follow-on innovation which may have spilled-over to products that use patented technologies in the dataset (hegde_patent_2018; kim_innovation_2021; baruffaldi_patents_2020; hegde_patent_2023). I control for this effect with a binary variable AIPA equal 1 for applications submitted by applicants with recent patents on or after its effective date of November 29, 2000.
A second set of changes the FCC implemented in 1998 changed the regulatory regime of several product categories, mainly by deregulating some products that do not actively emit electromagnetic signal (federal_communications_commission_fcc_1997; federal_communications_commission_fcc_1998). The deregulated products no longer had to acquire prior authorization from the FCC unless the firm submitted an application voluntarily. This implies that such products would stop appearing in the dataset. To avoid conflation from a selection bias stemming from the co-occurrence of the two events, I mapped regulatory changes to product categories and excluded all deregulated product from the analysis, including those submitted prior to the policy change. See Supplemental Appendix LABEL:apx:sec:deregulated for details and validation of the process.
The period of interest also saw the entry of the first wireless internet-connected devices. To control for the effects of this development, I map the frequencies of 3G, 4G and Wifi technologies, and code Internet as a binary variable equal 1 if the product had internet capability. Supplemental Appendix LABEL:apx:sec:internet lists these frequencies and details the process.
@@ 7.2 Foreign and domestic competition
Domestic $\times$ AIPA Yes Yes Yes Yes
@@ 8 Conclusion
Government regulations affect the information environment in which firms innovate. While the existence of such connection is undisputed, little empirical evidence exists on the effects of government policies that affect information provided by and available to private firms, especially outside of the patent system. To elucidate this dynamic, I evaluate the effects of increased regulatory transparency on follow-on innovation. I focus on a late-1990s transparency shock that decreased the costs of accessing detailed technical information about all wireless-enabled devices marketed in the U.S., a globally representative high-tech industry.
The results expose a familiar tension in innovation policy: eased access to information about a new technology benefits competitors and consumers but may hinder the original firm’s ability to offset the costs of developing the technology in the first place (hall_choice_2014). In the long term, this dynamic may discourage innovation. These considerations preclude an unequivocally positive evaluation of increased transparency’s overall effects on innovation, as such an evaluation necessitates modeling firms’ strategic responses within a dynamic equilibrium of the altered information environment (hall_appropriability_2017; bryan_innovation_2021).
The main message from my analysis is that government transparency can have large effects on business innovation in ways that are typically not at the forefront of policy debates on transparency. The findings highlight the information frictions present in innovation markets, and the power of government interventions in shaping the information environment in which firms innovate. Even regulations not directly aimed at supporting innovation can generate significant unintended effects by influencing information costs. As governments consider policies affecting information availability—such as algorithmic transparency, AI regulations, or restrictions on non-compete agreements—these results suggest potential secondary effects on innovation, even when this is not the primary objective. The FCC product database remains unaffected by recent discontinuations of U.S. government information resources, but numerous other government websites have been terminated (gotfredsen_fighting_2025). The results presented here show that such changes in government transparency may have broad repercussions.